OpenStudy (anonymous):

three people who work full time are to work together on a project, but their total time on the project is to be equivalant to that of only one person working full time. If one of the people is budgeted for one-half of his time to the project and a second person for one-third of her time, what part of the third workers time should be budgeted to this project????

OpenStudy (katy):

STEP 1: MAKE SENSE OF THE CRAZY WORDY MATHY STUFF. Breaking it down, you have person A, B, and C. All are working full time, but not all of the time spent working "full time" is being put towards the project. We'll just say that "full time"=x hours You have 1/2 of A's "x hours", 1/3 of B's, and ? of C's. The project has to equal one person's "full time." STEP 2: USE YOUR MATH-INATOR RAY TO TURN THE YUCKY WORD PROBLEM INTO A CONVENIENT LIST OF VARIABLES! (and variable relationships.) A=1/2x B=1/3x C=[unknown] x=1 A+B+C=x STEP 3: PLUG AND CHUG!! 1/2*(1)+1/3*(1)+C=(1) STEP 4: SOLVE FOR YOUR VARIABLE!!! STEP 5: THROW A PARTY-YOU FINISHED THE PROBLEM :D

OpenStudy (anonymous):

the answer should be one of the following ow would i determine this? A.13.3% B.35.2% C. 16.7% D.18.7%

OpenStudy (katy):

Your last step will land you with a fraction answer for C. Divide out the fraction to get a decimal, and multiply the decimal by 100 to get the percentage. Hope that helps!

OpenStudy (katy):

Are you good with finding the common denominators to combine the like terms earlier in the problem?